please solve the system
y=4x+3
4y-20x=-8

Answer by checkley71(8403)

Y=4X+3 THEN
4(4X+3)-20X=-8
16X+12-20X=-8
-4X=-8-12
-4X=-20
X=-20/-4
X=5
Y=4*5+3
Y=20+3
Y=23
PROOF
4*23-20*5=-8
92-100=-8
-8=-8